OpenStudy (anonymous):

I'll give medal!!! I need help on understanding how to factor quadratics. Here's the equation: (x^2-4x+4)+(y^2-2y+1)=9. I know the answer but I don't understand how to get there. Plz help:)

OpenStudy (mathstudent55):

When you have a trinomial that you need to factor, you need to check if it is a perfect square trinomial.

OpenStudy (anonymous):

How do I do that?

OpenStudy (mathstudent55):

Let's look at x^2 - 4x + 4 This may be (x - 2)^2, but we need to be sure.

OpenStudy (mathstudent55):

1) The first term must be a perfect square. If so, what is it the square of? 1) Answer: x^2 is the square of x.

OpenStudy (anonymous):

Ok.

OpenStudy (mathstudent55):

2) The last term must be a perfect square. If so, what is it the square of? 2) Answer: 4 is a perfect square, and it is the square of 2 and -2.

OpenStudy (mathstudent55):

That means this trinomial can possibly be (x + 2)^2 or (x - 2)^2. Now we need to check the middle term of the trinomial to see if it works.

OpenStudy (mathstudent55):

The middle term of a perfect square trinomial is in red below: \( (a + b)^2 = a^2 \color{red}{+ 2ab} + b^2\) \( (a - b)^2 = a^2 \color{red}{- 2ab} + b^2\)

OpenStudy (anonymous):

So for it to be perfect, it has to be 2 times the first and last term? Positive or negative.

OpenStudy (mathstudent55):

Exactly.

OpenStudy (mathstudent55):

Once you know what the first and last terms of the TRINOMIAL are squares of, the middle term of the trinomial is 2 times the first term times the second term of the BINOMIAL.

OpenStudy (anonymous):

So it would look like 2*(x*4)?

OpenStudy (mathstudent55):

Back to our trinomial. Since the middle term of the trinomial is negative, the only possibility is \((x - 2)^2\)

OpenStudy (mathstudent55):

Back to our trinomial. Since the middle term of the trinomial is negative, the only possibility is (x - 2)^2

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

So then the answer would be the square (or what ever the power is) of what we now know is the middle term?